6. Give the balanced full reaction and calculate the E° for galvanic cells based on the

following reactions. Draw diagrams indicating; the anode, cathode, flow of electrons, salt bridge

(a) Zn

2+

(aq)+ 2e

-  Zn (s)

Pb

2+

(aq) + 2e

-  Pb(s)

(b) Cr2O7

- + 14H+ + 6e

-  2Cr

3+ + 7H2O

H2O2 + 2H+ + 2e

-  2H2O


1
Expert's answer
2022-06-08T05:25:03-0400

a)

Zn2+ + 2 e- = Zn

Pb2+ + 2 e- = Pb

Full reaction: PbCl2 + Zn = ZnCl2 + Pb

Anode: Zn => Zn2+ + 2e-

Cathode: Pb2+ + 2e- => Pb

b)

Cr2O7- + 14H+ + 6 e- = 2 Cr3+ + 7H2O

H2O2 + 2H+ 2 e- = 2 H2O

Full reaction: K2Cr2O7 + 3 H2O2 + 4 H2SO4 = Cr2(SO4)3 + K2SO+ 7H2O + 3O2